# Generative Adversarial Networks for Image Super-Resolution: A Survey

### v4

@@ 4.3.2. Unsupervised GANs based prior knowledge for image super-resolution
Combining unsupervised GANs and prior knowledge in unsupervised GANs can better address unpair image super-resolution (Lin et al., 2018). Lin et al. combined data error, a regular term and an adversarial loss to guarantee consistency of local-global content and pixel faithfulness in a GAN in an unsupervised way to train an image super-resolution model (Lin et al., 2018). To better support medical diagnosis, Das et al. combined adversarial learning in a GAN, cycle consistency and prior knowledge, i.e., identity mapping prior to facilitate more useful information i.e., spatial correlation, color and texture information for obtaining cleaner high-quality images (Das et al., 2020). In terms of remoting sensing super-resolution, a random noise is used in a GAN to reconstruct satellite images (Wang et al., 2021b). Then, authors conducted image prior by transforming the reference image into a latent space (Wang et al., 2021b). Finally, they updated the noise and latent space to transfer obtained structure information and texture information for improving resolution of remote sensing images (Wang et al., 2021b).
@@ 4.3.3. Unsupervised GANs with improved loss functions for image super-resolution
EIPGAN (Wang et al., 2021b) GAN GAN with remote sensing image prior for image super-resolution
@@ 4.3.4. Unsupervised GANs based multi-tasks for image super-resolution
5. Comparing performance of GANs for image super-resolution
5.1. Datasets
(1) Supervised GANs for image-resolution
(2) Semi-supervised GANs for image-resolution
Test datasets: Tibia Dataset (Chen et al., 2018), Abdominal Dataset (McCollough et al., 2020), CUHK03 (Li et al., 2014), Widerface (Yang et al., 2016a), LUNA (Setio et al., 2017), DSB (Kuan et al., 2017), SG (Woodbridge et al., 2013) and UKBB (Bycroft et al., 2018).
(3) Unsupervised GANs for image-resolution
Training datasets: CIFAR10 (Krizhevsky et al., 2009), ImageNet (Russakovsky et al., 2015), DIV2K (Agustsson and Timofte, 2017), DIV2K random kernel (DIV2KRK) (Agustsson and Timofte, 2017), Flickr2K (Wang et al., 2019), Widerface (Yang et al., 2016a), NTIRE 2020 Real World SR challenge (Lugmayr et al., 2020), KADID-10K (Lin et al., 2019), DPED (Ignatov et al., 2017), DF2K (Wang et al., 2018b), NTIRE’ 2018 Blind-SR challenge (Timofte et al., 2018), LS3D-W (Bulat and Tzimiropoulos, 2017), CELEBA-HQ (Karras et al., 2017), LSUN-BEDROOM (Yu et al., 2015), ILSVRC2012 (Russakovsky et al., 2015; Nguyen et al., 2017), NTIRE 2020 (Lugmayr et al., 2020), 91-images (Yang et al., 2010), Berkeley segmentation (Martin et al., 2001), BSDS500 (Yang et al., 2016b), Training datasets of USROCTGAN (Fang et al., 2013, 2012), SD-OCT dataset (Fang et al., 2012), UC Merced dataset (Yang and Newsam, 2010), NWPU-RESIS45 (Cheng et al., 2017) and WHU-RS19 (Dai and Yang, 2010).
Test datasets: CIFAR10 (Krizhevsky et al., 2009), ImageNet (Russakovsky et al., 2015), Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2010), BSD100 (Martin et al., 2001), DIV2K (Agustsson and Timofte, 2017), DIV2KRK (Agustsson and Timofte, 2017), Urban100 (Huang et al., 2015), Widerface (Yang et al., 2016a), NTIRE 2020 (Lugmayr et al., 2020), NTIRE 2020 Real-world SR Challenge (Lugmayr et al., 2020), NTIRE 2020 Real World SR challenge validation dataset (Lugmayr et al., 2020), DPED (Ignatov et al., 2017), CELEBA-HQ (Karras et al., 2017), LSUN-BEDROOM (Yu et al., 2015), Test datasets of USROCTGAN (Fang et al., 2013, 2012), and Test datasets of USRGAN (Zhang et al., 2020).
@@ 5.1. Datasets
Training ways Training datasets Test datasets
Supervised ways
@@
Urban100 (Huang et al., 2015).
@@
PASCAL2 (Everingham et al., 2010).
Semi-supervised ways
@@
Market-1501 (Zheng et al., 2015),
CUHK03 (Li et al., 2014),
MSMT17 (Wei et al., 2018),
LUNA (Setio et al., 2017),
Data Science Bowl 2017 (DSB) (Kuan et al., 2017),
UKDHP (Woodbridge et al., 2013),
SG (Woodbridge et al., 2013),
UKBB (Bycroft et al., 2018).
Tibia Dataset (Chen et al., 2018),
Abdominal Dataset (McCollough et al., 2020),
Market-1501 (Zheng et al., 2015),
@@
UKBB (Bycroft et al., 2018).
Unsupervised ways
DIV2K (Agustsson and Timofte, 2017),
Flickr2K (Wang et al., 2019),
Widerface (Yang et al., 2016a),
NTIRE-2020 Real-world SR Challenge validation dataset (Lugmayr et al., 2020),
KADID-10K (Lin et al., 2019),
@@
DF2K (Wang et al., 2018b),
NTIRE’2018 Blind-SR challenge (Timofte et al., 2018),
DIV2K random kernel (DIV2KRK) (Agustsson and Timofte, 2017),
LS3D-W (Bulat and Tzimiropoulos, 2017),
CIFAR10 (Krizhevsky et al., 2009),
@@
NTIRE 2020 (Lugmayr et al., 2020),
91-images (Yang et al., 2010),
Berkeley segmentation (Martin et al., 2001), BSDS500 (Yang et al., 2016b),
Training datasets of USROCTGAN (Fang et al., 2013, 2012),
SD-OCT dataset (Fang et al., 2012),
UC Merced dataset (Yang and Newsam, 2010),
NWPU-RESIS45 (Cheng et al., 2017),
WHU-RS19 (Dai and Yang, 2010).
DIV2K (Agustsson and Timofte, 2017),
Set5 (Bevilacqua et al., 2012),
Set14 (Bevilacqua et al., 2012),
Urban100 (Huang et al., 2015),
BSD100 (Martin et al., 2001),
NTIRE 2020 Real World SR challenge (Lugmayr et al., 2020),
NTIRE-2020 Real-world SR Challenge validation dataset (Lugmayr et al., 2020),
DPED (Ignatov et al., 2017),
DIV2KRK (Agustsson and Timofte, 2017),
Widerface (Yang et al., 2016a),
CIFAR10 (Krizhevsky et al., 2009),
ImageNet (Russakovsky et al., 2015),
CELEBA-HQ (Karras et al., 2017),
LSUN-BEDROOM (Yu et al., 2015),
NTIRE 2020 (Lugmayr et al., 2020),
@@
DIV2K (Agustsson and Timofte, 2017), Flickr2K (Wang et al., 2019), Widerface (Yang et al., 2016a)
@@
NTIRE 2020 Real World SR challenge (Lugmayr et al., 2020)
NTIRE-2020 Real-world SR Challenge validation dataset (Lugmayr et al., 2020), DIV2K (Agustsson and Timofte, 2017), Flickr2k (Wang et al., 2019), KADID-10K (Lin et al., 2019)
@@
DIV2K (Agustsson and Timofte, 2017), NTIRE 2020 (Lugmayr et al., 2020)
91-images (Yang et al., 2010), Berkeley segmentation (Martin et al., 2001), BSDS500 (Yang et al., 2016b)
@@
NTIRE 2020 Real World SR challenge (Lugmayr et al., 2020)
NTIRE-2020 Real-world SR Challenge validation dataset (Lugmayr et al., 2020)
@@
DIV2K (Agustsson and Timofte, 2017), NTIRE 2020 (Lugmayr et al., 2020)
@@ 5.3.1. Quantitative analysis of different GANs for image super-resolution
ESRGAN (Wang et al., 2018b) 9.7607 1.670×$10^{7}$
PathSRGAN (Ma et al., 2020) 9.1608 2.433×$10^{7}$
RankSRGAN (Zhang et al., 2019b) 7.3818 1.554×$10^{6}$
KernelGAN (Bell-Kligler et al., 2019) 251.00 1.816×$10^{5}$
@@ 5.3.2. Qualitative analysis of different GANs for image super-resolution
To test visual effects of different GANs for image super-resolution, we choose Bicubic, ESRGAN (Wang et al., 2018b), RankSRGAN (Zhang et al., 2019b), KernelGAN (Bell-Kligler et al., 2019) and PathSRGAN (Ma et al., 2020) to conduct experiments to obtain high-quality images for ×4. To further observe these images, we choose an area of predicted images from these GANs to amplify it as an observation area. Observation area is clearer, corresponding method has good superior SR performance. For example, ESRGAN (Wang et al., 2018b) is clearer than that of PathSRGAN (Ma et al., 2020) on an image from the BSD100 in Fig. 9 and Set14 in Fig. 10 for ×4, which show that the ESRGAN is more effective in image super-resolution.
[…]
Qualitative experiment on BSD100.
[…]
Qualitative experiment on Set14.
@@ 7. Conclusion
In this paper, we analyze and summarize GANs for image super-resolution. First, we introduce developments of GANs. Then, we present GANs in big samples and small samples for image applications, which can make readers easier GANs. Next, we give differences of GANs based optimization methods and discriminative learning for image super-resolution in terms of supervised, semi-supervised and unsupervised manners. Subsequently, we compare the performance of these popular GANs on public datasets via quantitative and qualitative analysis in SISR. Finally, we highlight challenges of GANs and potential research points on SISR.
@@ Acknowledgements.
This work was supported in part by National Natural Science Foundation of China under Grant 62201468, in part by the China Postdoctoral Science Foundation Grant 2022TQ0259, in part by the Jiangsu Provincial Double–Innovation Doctor Program Grant JSSCBC20220942. The authors appreciate informative and insightful comments provided by anonymous editors and reviewers, which significantly improve the quality of this survey.

### v5

@@ 4.3.2. Unsupervised GANs based prior knowledge for image super-resolution
Combining unsupervised GANs and prior knowledge in unsupervised GANs can better address unpair image super-resolution (Lin et al., 2018). Lin et al. combined data error, a regular term and an adversarial loss to guarantee consistency of local-global content and pixel faithfulness in a GAN in an unsupervised way to train an image super-resolution model (Lin et al., 2018). To better support medical diagnosis, Das et al. combined adversarial learning in a GAN, cycle consistency and prior knowledge, i.e., identity mapping prior to facilitate more useful information i.e., spatial correlation, color and texture information for obtaining cleaner high-quality images (Das et al., 2020). In terms of remoting sensing super-resolution, a random noise is used in a GAN to reconstruct satellite images (Wang et al., 2021a). Then, authors conducted image prior by transforming the reference image into a latent space (Wang et al., 2021a). Finally, they updated the noise and latent space to transfer obtained structure information and texture information for improving resolution of remote sensing images (Wang et al., 2021a).
@@ 4.3.3. Unsupervised GANs with improved loss functions for image super-resolution
EIPGAN (Wang et al., 2021a) GAN GAN with remote sensing image prior for image super-resolution
@@ 4.3.4. Unsupervised GANs based multi-tasks for image super-resolution
4.3.5. Unsupervised GANs for real image super-resolution
An important application of GAN-based SR models is dealing with real world LR images. To eliminate checkerboard artifacts, an upsampling module containing a bilinear interpolation and a transposed convolution was used in an unsupervised CycleGAN to improve visual effects of restored images in the real world (Kim et al., 2020). To recover more natural image characteristics, Lugmayr et al. combined unsupervised and supervised ways for blind image super-resolution (Lugmayr et al., 2019). The first step learned to invert the effects of bicubic down sampling operation in a GAN in an unsupervised way to extract useful information from natural images (Lugmayr et al., 2019). To generate image pairs in the real world, the second step used a pixel-wise network in a supervised way to obtain high-resolution images (Lugmayr et al., 2019). To remove unpleasing noise and artifacts, Park et al. (Park et al., 2023b) divided the large and complicated distribution of real-world images into smaller subsets based on similar contents. Then, they learned the various contents and separable features via multiple discriminators (Park et al., 2023b). In terms of computed tomography (CT) images, Li et al. (Li et al., 2023) utilized noise injection and a GAN to generate realistic low-resolution images to build training image pairs. This data preprocessing operation enables the proposed KerGAN to generate more precise details and better perceptual quality medical images (Li et al., 2023).
4.3.6. Synthesizing real world LR images
An important application of GAN-based SR models is dealing with real world LR images. So, how to obtain real image pairs with different scale factors to be used as training data becomes a problem that we must solve. Starting from SRCNN (Dong et al., 2015), there are three main ways regarding the preparation of super-resolution datasets: synthetic data super resolution, blind super-resolution and real image super-resolution. Among them, synthetic data super-resolution is the most common way of data preparation, and this route usually uses the intuitive interpolation method as a degradation kernel for downsampling to obtain the corresponding training image pairs. However, in the face of complex degradation factors in the real world, synthetic data often do not generalize well. Therefore, on the basis of the interpolation method, blind super-resolution models add more real world degradation factors, i.e., noise, blur, compression etc. to the degradation kernel to improve the performance of real world applications (Wang et al., 2021b; Ren et al., 2020). Finally, the best theoretical performance is achieved by directly utilizing a camera to capture real world images of the same scene at different magnifications. However, in this process, the problem of lens distortion is encountered which makes the image acquisition difficult. Currently, the publicly available real image datasets are RealSR (Cai et al., 2019) and DRealSR (Wei et al., 2020).
@@ 4.3.6. Synthesizing real world LR images
Training datasets Test datasets
@@
Urban100 (Huang et al., 2015),
Tibia Dataset (Chen et al., 2018),
Abdominal Dataset (McCollough et al., 2020),
CUHK03 (Li et al., 2014),
MSMT17 (Wei et al., 2018),
LUNA (Setio et al., 2017),
Data Science Bowl 2017 (DSB) (Kuan et al., 2017),
UKDHP (Woodbridge et al., 2013),
SG (Woodbridge et al., 2013),
UKBB (Bycroft et al., 2018),
Widerface (Yang et al., 2016b),
NTIRE-2020 Real-world SR Challenge (Lugmayr et al., 2020a),
KADID-10K (Lin et al., 2019),
DPED (Ignatov et al., 2017),
DF2K (Wang et al., 2018b),
NTIRE’2018 Blind-SR challenge (Timofte et al., 2018),
DIV2K random kernel (DIV2KRK) (Agustsson and Timofte, 2017),
LS3D-W (Bulat and Tzimiropoulos, 2017),
CELEBA-HQ (Karras et al., 2017),
LSUN-BEDROOM (Yu et al., 2015),
NTIRE 2020 (Lugmayr et al., 2020a),
91-images (Yang et al., 2010),
Berkeley segmentation (Martin et al., 2001),
BSDS500 (Yang et al., 2016a),
Training datasets of USROCTGAN (Fang et al., 2013, 2012),
SD-OCT dataset (Fang et al., 2012),
UC Merced dataset (Yang and Newsam, 2010),
NWPU-RESIS45 (Cheng et al., 2017),
WHU-RS19 (Dai and Yang, 2010).
@@
PASCAL2 (Everingham et al., 2010),
@@
UKBB (Bycroft et al., 2018),
Urban100 (Huang et al., 2015),
NTIRE 2020 Real World SR challenge (Lugmayr et al., 2020a),
@@
DIV2KRK (Agustsson and Timofte, 2017),
Widerface (Yang et al., 2016b),
@@
5. Comparing performance of GANs for image super-resolution
5.1. Datasets
(1) Supervised GANs for image-resolution
(2) Semi-supervised GANs for image-resolution
Test datasets: Tibia Dataset (Chen et al., 2018), Abdominal Dataset (McCollough et al., 2020), CUHK03 (Li et al., 2014), Widerface (Yang et al., 2016b), LUNA (Setio et al., 2017), DSB (Kuan et al., 2017), SG (Woodbridge et al., 2013) and UKBB (Bycroft et al., 2018).
(3) Unsupervised GANs for image-resolution
Training datasets: CIFAR10 (Krizhevsky et al., 2009), ImageNet (Russakovsky et al., 2015), DIV2K (Agustsson and Timofte, 2017), DIV2K random kernel (DIV2KRK) (Agustsson and Timofte, 2017), Flickr2K (Wang et al., 2019), Widerface (Yang et al., 2016b), NTIRE 2020 Real World SR challenge (Lugmayr et al., 2020a), KADID-10K (Lin et al., 2019), DPED (Ignatov et al., 2017), DF2K (Wang et al., 2018b), NTIRE’ 2018 Blind-SR challenge (Timofte et al., 2018), LS3D-W (Bulat and Tzimiropoulos, 2017), CELEBA-HQ (Karras et al., 2017), LSUN-BEDROOM (Yu et al., 2015), ILSVRC2012 (Russakovsky et al., 2015; Nguyen et al., 2017), NTIRE 2020 (Lugmayr et al., 2020a), 91-images (Yang et al., 2010), Berkeley segmentation (Martin et al., 2001), BSDS500 (Yang et al., 2016a), Training datasets of USROCTGAN (Fang et al., 2013, 2012), SD-OCT dataset (Fang et al., 2012), UC Merced dataset (Yang and Newsam, 2010), NWPU-RESIS45 (Cheng et al., 2017) and WHU-RS19 (Dai and Yang, 2010).
Test datasets: CIFAR10 (Krizhevsky et al., 2009), ImageNet (Russakovsky et al., 2015), Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2010), BSD100 (Martin et al., 2001), DIV2K (Agustsson and Timofte, 2017), DIV2KRK (Agustsson and Timofte, 2017), Urban100 (Huang et al., 2015), Widerface (Yang et al., 2016b), NTIRE 2020 (Lugmayr et al., 2020a), NTIRE 2020 Real-world SR Challenge (Lugmayr et al., 2020a), NTIRE 2020 Real World SR challenge validation dataset (Lugmayr et al., 2020a), DPED (Ignatov et al., 2017), CELEBA-HQ (Karras et al., 2017), LSUN-BEDROOM (Yu et al., 2015), Test datasets of USROCTGAN (Fang et al., 2013, 2012), and Test datasets of USRGAN (Zhang et al., 2020).
@@ 5.1. Datasets
DIV2K (Agustsson and Timofte, 2017), Flickr2K (Wang et al., 2019), Widerface (Yang et al., 2016b)
@@
NTIRE 2020 Real World SR challenge (Lugmayr et al., 2020a)
NTIRE-2020 Real-world SR Challenge validation dataset (Lugmayr et al., 2020a), DIV2K (Agustsson and Timofte, 2017), Flickr2k (Wang et al., 2019), KADID-10K (Lin et al., 2019)
@@
DIV2K (Agustsson and Timofte, 2017), NTIRE 2020 (Lugmayr et al., 2020a)
91-images (Yang et al., 2010), Berkeley segmentation (Martin et al., 2001), BSDS500 (Yang et al., 2016a)
@@
NTIRE 2020 Real World SR challenge (Lugmayr et al., 2020a)
NTIRE-2020 Real-world SR Challenge validation dataset (Lugmayr et al., 2020a)
@@
DIV2K (Agustsson and Timofte, 2017), NTIRE 2020 (Lugmayr et al., 2020a)
@@ 5.3.1. Quantitative analysis of different GANs for image super-resolution
[…]
@@ 5.3.2. Qualitative analysis of different GANs for image super-resolution
To test visual effects of different GANs for image super-resolution, we choose Bicubic, ESRGAN (Wang et al., 2018b), RankSRGAN (Zhang et al., 2019b), KernelGAN (Bell-Kligler et al., 2019) and PathSRGAN (Ma et al., 2020) to conduct experiments to obtain high-quality images for ×4. To further observe these images, we choose an area of predicted images from these GANs to amplify it as an observation area. Observation area is clearer, corresponding method has good superior SR performance. For example, ESRGAN (Wang et al., 2018b) is clearer than that of PathSRGAN (Ma et al., 2020) on an image from the BSD100 in Fig 9 and Set14 in Fig 10 for ×4, which show that the ESRGAN is more effective in image super-resolution.
@@ 6. Challenges and directions of GANs for image super-resolution
6) GANs are sensitive to hyperparameters and the choice of architecture, which can affect their performance and lead to overfitting.
7) GANs can generate unrealistic artifacts or distortions in the output images, especially in regions with low texture or detail, which can affect the visual quality of the output images.
8) GAN-based image super-resolution methods may not perform well on certain types of images or degradation models.
@@ 7. Conclusion
In this paper, we analyze and summarize GANs for image super-resolution. First, we review the development of GANs and popular GANs for image applications; then we give differences of GANs based optimization methods and discriminative learning for image super-resolution in terms of supervised, semi-supervised and unsupervised manners. Next, we compare the performance of these popular GANs on public datasets via quantitative and qualitative analysis in SISR. Then, we highlight challenges of GANs and potential research points on SISR. Finally, we summarize the whole paper.
@@ Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China under Grants 62576123, in part by CAAI-CANN Open Fund developed on OpenI Community, and in part by the Natural Science Foundation of Heilongjiang Province under Grant YQ2025F003.